# DeepCRE: Revolutionizing Drug R&D with Cutting-Edge Computational Models

### 09 Mar 2024 v2

@@ Abstract
The fields of pharmaceutical development and therapeutic application both face substantial challenges. The therapeutic domain calls for more treatment alternatives, while numerous promising pre-clinical drugs have failed in clinical trials. One of the reasons is the inadequacy of Cross-drug Response Evaluation (CRE) during the late stage of drug development. Although in-silico CRE models offer a solution to this problem, existing methodologies are either limited to early development stages or lack the capacity for a comprehensive CRE analysis. Herein, we introduce a novel computational model named DeepCRE and present the potential of DeepCRE in advancing therapeutic discovery and development. DeepCRE outperforms the existing best models by achieving an average performance improvement of 17.7% in patient-level CRE and a 5-fold increase in indication-level CRE. Furthermore, DeepCRE has identified six drug candidates that show significantly greater effectiveness than a comparator set of two approved drugs in 5/8 colorectal cancer (CRC) organoids. This highlights DeepCRE’s ability to identify a collection of drug candidates with superior therapeutic effects, underscoring its potential to revolutionize the field of therapeutic development.
@@ Introduction
The fields of pharmaceutical development and therapeutic application both face substantial challenges[1, 2, 3, 4]. In therapeutic applications, the scarcity of effective treatment options is a critical concern for medical researchers[2]. This challenge was highlighted in a recent article in Nature, entitled ”The future of precision cancer therapy might be to try everything”, advocating for a comprehensive exploration of treatment alternatives[1]. Unfortunately, the reality is characterized by a constricted repertoire of available drugs[2]. On the other hand, the domain of drug development confronts the phenomenon known as ”Eroom’s Law”, which posits that, counterintuitively, the efficiency of developing new drugs is declining despite advancements in technology[5]. The failure of numerous promising pre-clinical drugs during clinical trials highlights the obstacles to expanding the spectrum of effective treatments.
@@
A pivotal factor exacerbating these challenges is the inadequacy of Cross-drug Response Evaluation (CRE), which is essential for reducing costs and enhancing efficiency in the drug development stage[6], and for improving prognosis outcomes in the therapeutic stage[2, 7, 8]. However, due to technological and ethical constraints, CRE has largely been restricted to the target and cell-line level[9, 10], representing the early stage of drug development (Fig. 1a). Hopefully, Artificial Intelligence (AI) has emerged as a key technology with the potential to predict the CRE in the late stage (Fig. 1a), offering new approach to exploring treatment strategies more effectively[11, 12].
@@
Recently, there have been some related researches [13, 14, 10, 15, 16]. Single-Drug Learning (SDL) models, such as ADAE[13] and Code-AE[14], have demonstrated potential in developing personalized treatment and identifying patient biomarkers. Nonetheless, these models fall short in the capability of comprehensive CRE, restricting their broader applications. In contrast, Multi-Drug Learning (MDL) models, including DrugCell[10], Paccmann[15], and TGSA[16], leverage extensive drug response data and facilitate cross-drug response comparisons [17]. However, MDL models are generally restricted to the cell-line level [17], which may result in suboptimal performance when predicting patient drug responses due to the disparities between cell lines and patients.
Here, we introduce DeepCRE, a novel computational model that significantly enhances CRE at the patient level, demonstrating an average performance boost of 17.77% over existing state-of-the-art (SOTA) models[13, 14, 10, 15, 16]. Furthermore, our model shows a remarkable 5-fold improvement in indication-level CRE, outperforming all previous SOTA models. Finally, we present compelling evidence of DeepCRE’s potential in drug repurposing for colorectal cancer (CRC). The model has identified a set of six drug candidates (Set A) that are significantly more effective than a comparator set of two approved drugs (Set C) in 5/8 CRC organoids testing, underscoring the potential of DeepCRE in advancing therapeutic discovery and development. To note, it is not merely about identifying one or two potential drugs, rather, it highlights the capability of DeepCRE to uncover a collection of drug candidates with enhanced therapeutic effects.
@@ Results
The drug research and design (R&D) process encompasses target identification, cell line assays, animal or organoid testing, and clinical trials [18]. A significant number of drug candidates fail at various stages, resulting in inefficient drug development[5, 19]. Interestingly, there is a magical connection between the success rate and the CRE abundance in every stage of drug R&D, where multiple CRE exist in the early stage while few CRE remains in the late stage (Fig. 1a). Enhancing the presence of CREs in later stages, especially with the aid of in-silico CRE models, could markedly improve drug development outcomes (Fig. 1a,b).
The principle of DeepCRE is to project the cell line and patient data onto a common space, which is suitable for drug response prediction (Fig. 1c). This methodology has contributed to an average performance boost of 17.7% and a 5-fold improvement in patient and indication-level CRE prediction, respectively, surpassing existing SOTA models (Fig. 1c). Notably, DeepCRE has identified six drug candidates (Set A) that exhibit significantly higher effectiveness than two approved drugs (Set C) in 5 out of 8 CRC organoids tested (Fig. 1c-e). To note, it is not merely about identifying one or two potential drugs, rather, it highlights the capability of DeepCRE to uncover a collection of drug candidates with enhanced therapeutic effects.
@@
We constructed the first patient-level CRE datasets and established the initial DeepCRE model (Supplementary Fig. 1-3 and Supplementary Table 1). Upon identifying limitations the limitations of the initial DeepCRE model (details refer to Supplementary materials, Supplementary Fig. 4a,5), we turned to the architecture of domain separation network (DSN), which is a strategy for transfer learning and has been successfully applied in computer vision[20, 21, 22]. We propose the DeepCRE model zoo, comprising a suite of the DSN architecture (Fig. 2a) and its derivatives (Fig. 2b). The DeepCRE model zoo achieves an average performance increase of 17.21%, 11.52% and 7.69% over P-SDL models, C-MDL models and their combined counterparts, respectively (Supplementary Fig. 4b). Nonetheless, there remain four tumor types in which the DeepCRE model zoo demonstrates only average performance (Supplementary Fig. 5). Additionally, we observe no distinct differences in performance among these DeepCRE models (Supplementary Table 2). Since these models primarily differ in pretraining stages, we hypothesize that some pretraining settings may not be suitable for specific tumor types.
@@
To further validate the predictions, we cross-reference in-clinical-test records from the ClinicalTrials database (Supplementary Table 11). Drug candidates identified by DeepCRE exhibit anti-tumor evidence in terms of mechanism of action (MoA) and current indications (Supplementary Fig. 10a,b). Nine drug candidates are in clinical testing, and after excluding widely-used Gemcitabine (Supplementary Fig. 10c,d), they are verified to be tested in various indications, including the 13 types of tumor indications used in our study (Fig. 6f and Supplementary Table 12). Mirroring the DrugBank analysis, the DEI table is visualized using a heatmap, which also includes additional information about the tumor types in which the drug candidates were undergoing clinical testing (Fig. 6g). The predicted results demonstrate excellent agreement with the in-clinical-test records, indicating strong concordance between DeepCRE predictions and real-world clinical data (details refer to the Supplementary materials). Furthermore, five drug candidates are identified as Qualified drugs by DeepCRE (Fig. 6g), while no more than one Qualified drug candidate is discovered by any previous SOTA models (Supplementary Fig. 11). Additionally, the EDIs and Qualified drugs identified by our model are associated with 115 and 105 in-clinical-test records, respectively, representing substantial increases of 117% and 400%, respectively, compared to previous SOTA models (Fig. 6h,i). This result means the drug candidates discovered by our model show at least a 5-fold increase in indication-level CRE compared to any previous SOTA models.
@@
DeepCRE’s capability to identify efficient drug candidates is further validated in a clinical setting, focusing on the CRC21 patient. Despite receiving XEOLX treatment, the patient experienced relapse after 13 days. Tumor samples collected before treatment are used to construct CRC organoids, enabling the evaluation of drug candidates identified by DeepCRE (Fig. 5a). This In-SMARchip drug testing methodology has been previously validated in lung and colorectal tumor organoids[30, 31, 32]. The drug candidates are categorized into four sets based on their MoAs and the evaluation variance between DeepCRE and traditional methods (Fig. 5b,c). We try to keep the MoAs diversity in each drug set (Fig. 5b) and ensure the drug set with the largest evaluation disparity for comparison (Fig. 5c).
@@
To further validate the potential of DeepCRE in drug repurposing for CRC, seven more CRC organoids are included for In-SMARchip drug testing. Among the eight CRC organoids tested, five exhibits a significant efficacy increase of set A over set C (Fig. 1d,e and Fig. 6a). Moreover, drug candidates in set A are also more efficient than two approved drugs in set C in the other three CRC organoids (Supplementary Fig. 13). Generally, the efficacy patterns of drug candidates were consistent across different patients, except for Sorafenib (D3), one of the receptor tyrosine kinase (RTK) inhibitors (Fig. 6b). However, there were variations in the exact correlation coefficients: inhibitors of PI3K/mTOR signaling and Chromatin histone acetylation show coefficients above 0.96, while those targeting DNA replication display coefficients ranging from 0.33 to 0.93, reflecting the diversity and complexity of chemotherapy[33, 34] (Fig. 3c).
@@
Some intriguing insights emerge from the drug testing results, shedding light on the deeper mechanism exploration of drug candidates with similar MoAs. For instance, the DRC of GSK1059615 exhibits a significant increase at the highest concentration compared to other PI3K/mTOR inhibitors (Fig. 6d). Perturbation data from CMAP L1000[35] database provides some explanations (Fig. 6e-f). Cells treated with GSK1059615 show up-regulated cellular survival and stress response related pathways (isomerase activity, intramolecular transferase activity, and phosphotransferases) and genes (IGFBP3, MYCBP, GTF2A2), DNA repair, and cytoskeleton regulation related genes (PARP1, RUVBL1, STAMBP), and down-regulated cell cycle checkpoints related pathways (protein tyrosine/MAP kinase/MAP kinase tyrosine phosphatase activity) and cell cycle related genes (STX1A, PARP6, KDM3A). The up-regulation and down-regulation of these genes and pathways may reflect the defensive mechanisms adopted by cells to resist the toxicity of high concentrations of drugs, potentially leading to the pausing of the cell cycle or the inhibition of cell death pathways[36, 37]. The other example is that the DRC of the drug Sorafenib, rather than another RTK inhibitor Dasatinib, declines at higher concentrations in three CRC organoids (Fig. 6g). It has been reported that some transport proteins may play a role in the accumulation and resistance of drugs, with Dasatinib being primarily restricted by ABCG1, while Sorafenib being mainly limited by ABCG2[38]. Further investigation suggests that ABCG2 expression, significantly lower in CRC tumor tissues compared to normal tissues[39] (Fig. 6h), may contribute to the reduced resistance of Sorafenib in CRC. Molecular docking analysis corroborates these findings, demonstrating a favorable binding affinity of Sorafenib to the target ABCG2 (Fig. 6i). Similar conclusions of binding poses and interactions have been drawn by a published paper, which presents single-particle cryo-EM studies of ABCG2 binding with another RTK inhibitor imatinib[40]. Furthermore, survival analysis reveals a negative correlation between ABCG2 expression and disease-free survival (DFS), implicating ABCG2-mediated multi-drug resistance (MDR) in CRC (Fig. 6j).
@@
a, DRC across four CRC organoids, colored by the drug set, indicating that drug set A possesses significantly greater effectiveness than drug set C. b, Heatmap of correlation coefficients of drugs, demonstrating that the drug efficacy is related to the drug MoA and the drug set. c, Boxplot of correlation coefficients of drugs, indicating differences in the distribution of efficient populations among drugs targeting DNA replication. d, DRC of PI3K/mTOR inhibitor drugs across four CRC organoids, showing the dramatic increase in DRC of the drug GSK1059615 at the highest concentration. e, Umap plot of gene expression of cells after different perturbations, with representative differential genes shown in the dashed frame. f, Bar plot of enriched pathways, with up-regulated pathways colored in red and down-regulated pathways in blue. g, DRC of RTK inhibitor drugs across three CRC organoids, showing the decline in DRC of the drug Dasatinib, rather than Sorafenib, at higher concentrations. h, Boxplot of genes ABCG1 and ABCG2 in the TCGA-CRC cohort, indicating significantly lower expression of ABCG2 in tumor samples. i, 3D and 2D diagram of ABCG2-Sorafenib docking. j, Survival curve of ABCG2 in the disease-free survival (DFS) of the TCGA-CRC cohort.
*** DISCUSSION ***

### 18 Mar 2024 v3

# DeepCRE: Transforming Drug R&D via AI-Driven Cross-drug Response Evaluation

@@ Abstract
The fields of therapeutic application and drug research and development (R&D) both face substantial challenges, i.e., the therapeutic domain calls for more treatment alternatives, while numerous promising pre-clinical drugs have failed in clinical trials. One of the reasons is the inadequacy of Cross-drug Response Evaluation (CRE) during the late stages of drug R&D. Although in-silico CRE models bring a promising solution, existing methodologies are restricted to early stages of drug R&D, such as target and cell-line levels, offering limited improvement to clinical success rates. Herein, we introduce DeepCRE, a pioneering AI model designed to predict CRE effectively in the late stages of drug R&D. DeepCRE outperforms the existing best models by achieving an average performance improvement of 17.7% in patient-level CRE, and a 5-fold increase in indication-level CRE, facilitating more accurate personalized treatment predictions and better pharmaceutical value assessment for indications, respectively. Furthermore, DeepCRE has identified a set of six drug candidates that show significantly greater effectiveness than a comparator set of two approved drugs in 5/8 colorectal cancer organoids. This demonstrates the capability of DeepCRE to systematically uncover a spectrum of drug candidates with enhanced therapeutic effects, highlighting its potential to transform drug R&D.
@@ Introduction
The fields of therapeutic application and drug research and development (R&D) both face substantial challenges [1, 2, 3, 4]. On the one hand, in therapeutic applications, the limited efficacy of existing treatment options is a critical concern for medical researchers [2]. This challenge is highlighted in a recent article in Nature, entitled ”The future of precision cancer therapy might be to try everything”, advocating for a comprehensive exploration of treatment alternatives [1]. Unfortunately, the reality is characterized by a constricted repertoire of available drugs [2]. On the other hand, the domain of drug R&D confronts the phenomenon known as ”Eroom’s Law”, which posits that, counterintuitively, the efficiency of developing new drugs is declining despite advancements in technology [5]. The failure of numerous promising pre-clinical drugs during clinical trials highlights the obstacles to expanding the spectrum of effective treatments.
@@
A pivotal factor exacerbating these challenges is the inadequacy of Cross-drug Response Evaluation (CRE). CRE enables the identification of the most effective drugs at various stages of drug R&D, which contributes to reducing costs and improving efficiency during the drug development stage [6], as well as in improving prognosis outcomes during the therapeutic stage [2, 7, 8]. However, due to technological and ethical constraints, CRE has largely been restricted to the target and cell line levels [9, 10], which represent the early stage of drug development (Fig. 1a). Hopefully, Artificial Intelligence (AI) has emerged as a key technology with the potential to predict the CRE in the late stage (Fig. 1a), offering a new approach to exploring treatment strategies more effectively [11, 12].
@@
Recently, a line of research have been proposed to develop in-silico CRE models [13, 14, 10, 15, 16], while they all come with some limitations. Single-Drug Learning (SDL) models, such as ADAE [13] and Code-AE [14], have demonstrated potential in developing personalized treatment and identifying patient biomarkers. Nonetheless, the SDL paradigm, employing a independent model for each specific drug, lacks CRE capabilities due to difficulties in comparing results across different models (Supplementary Fig. 1a). This limitation hampers their ability to identify the most effective drug for specific patients and indications, thereby restricting their broader applications. In contrast, Multi-Drug Learning (MDL) models, including DrugCell [10], Paccmann [15], and TGSA [16], leverage extensive drug response data and facilitate cross-drug response comparisons [17]. However, these MDL models are generally restricted to the cell-line level [17], which may result in suboptimal performance when predicting patient-level drug responses due to the disparities between cell lines and patients.
In this paper, we introduce DeepCRE, a pioneering AI model designed to predict CRE effectively in the late stages of drug R&D. This approach aims to tackle the substantial challenge of the limited effective treatment options in both therapeutic applications and drug R&D. We demonstrate the utility of DeepCRE through patient- and indication-level CRE analyses, highlighting its unique advantages for therapeutic applications. To further substantiate the transforming potential of DeepCRE in drug R&D, we employ wet lab assays, which validate its superior performance. Specifically, DeepCRE outperforms the existing state-of-the-art (SOTA) models [13, 14, 10, 15, 16] by achieving an average performance improvement of 17.7% in patient-level CRE, and a 5-fold increase in indication-level CRE. These advancements enable more precise predictions of personalized treatments and offer improved pharmaceutical value assessment for specific indications, respectively. Furthermore, DeepCRE has identified a set of six drug candidates that show significantly greater effectiveness than a comparator set of two approved drugs in 5/8 colorectal cancer organoids. To note, it is not merely about identifying one or two potential drugs, rather, it highlights the capability of DeepCRE to systematically uncover a spectrum of drug candidates with enhanced therapeutic effects.
@@ Results
The drug research and design (R&D) process encompasses target identification, cell line assays, animal or organoid testing, and clinical trials [18]. A significant number of drug candidates fail at various stages, resulting in inefficient drug development [5, 19]. Interestingly, there is a magical connection between the success rate and the CRE abundance in every stage of the drug R&D, where multiple CRE exist in the early stage while few CRE remains in the late stage (Fig. 1a). Enhancing the presence of CREs in later stages, especially with the aid of in-silico CRE models, could markedly improve drug development outcomes (Fig. 1a,b).
The principle of DeepCRE is to project the cell line and patient data onto a shared space (Fig. 1c). This approach enables the use of a drug response prediction model, trained using cell line data, to accurately predict patient drug responses. This methodology has contributed to an average performance boost of 17.7% and a 5-fold improvement in patient and indication-level CRE prediction, respectively, surpassing existing SOTA models (Fig. 1c). Notably, DeepCRE has identified six drug candidates (Set A) that exhibit significantly higher effectiveness than two approved drugs (Set C) in 5 out of 8 CRC organoids tested (Fig. 1c-e). To note, it is not merely about identifying one or two potential drugs, rather, it highlights the capability of DeepCRE to uncover a collection of drug candidates with enhanced therapeutic effects.
@@
We constructed the first patient-level CRE dataset and established the initial DeepCRE model (Supplementary Fig. 1-3 and Supplementary Table 1). Upon identifying the limitations of the initial DeepCRE model (details refer to Supplementary materials, Supplementary Fig. 4a,5), we turned to the architecture of domain separation network (DSN), which is a strategy for transfer learning and has been successfully applied in computer vision [20, 21, 22]. We propose the DeepCRE model zoo, comprising a suite of the DSN architecture (Fig. 2a) and its derivatives (Fig. 2b). The DeepCRE model zoo achieves an average performance increase of 17.21%, 11.52% and 7.69% over P-SDL models, C-MDL models and their combined counterparts, respectively (Supplementary Fig. 4b). Nonetheless, there remain four tumor types in which the DeepCRE model zoo demonstrates only average performance (Supplementary Fig. 5). Additionally, we observe no distinct differences in performance among these DeepCRE models (Supplementary Table 2). Since these models primarily differ in pretraining stages, we hypothesize that some pretraining settings may not be suitable for specific tumor types.
@@
To further validate the predictions, we cross-reference in-clinical-test records from the ClinicalTrials database (Supplementary Table 11). Drug candidates identified by DeepCRE exhibit anti-tumor evidence in terms of mechanism of action (MoA) and current indications (Supplementary Fig. 10a,b). Nine drug candidates are in clinical testing, and after excluding widely-used Gemcitabine (Supplementary Fig. 10c,d), they are verified to be tested in various indications, including the 13 types of tumor indications used in our study (Fig. 6f and Supplementary Table 12). Mirroring the DrugBank analysis, the DEI table is visualized using a heatmap, which also includes additional information about the tumor types in which the drug candidates were undergoing clinical testing (Fig. 6g). The predicted results demonstrate excellent agreement with the in-clinical-test records, indicating strong concordance between DeepCRE’s predictions and real-world clinical data (details refer to the Supplementary materials). Furthermore, five drug candidates are identified as Qualified drugs by DeepCRE (Fig. 6g), while no more than one Qualified drug candidate is discovered by any previous SOTA models (Supplementary Fig. 11). Additionally, the EDIs and Qualified drugs identified by our model are associated with 115 and 105 in-clinical-test records, respectively, representing substantial increases of 117% and 400%, respectively, compared to previous SOTA models (Fig. 6h,i). This result means the drug candidates discovered by our model show at least a 5-fold increase in indication-level CRE compared to any previous SOTA models.
@@
DeepCRE’s capability to identify efficient drug candidates is further validated in a clinical setting, focusing on the CRC21 patient. Despite receiving XEOLX treatment, the patient experienced relapse after 13 days. Tumor samples collected before treatment are used to construct CRC organoids, enabling the evaluation of drug candidates identified by DeepCRE (Fig. 5a). This In-SMARchip drug testing methodology has been previously validated in lung and colorectal tumor organoids [30, 31, 32]. The drug candidates are categorized into four sets based on their MoAs and the evaluation variance between DeepCRE and traditional methods (Fig. 5b,c and Supplementary Fig. 13). We try to keep the MoAs diversity in each drug set (Fig. 5b) and ensure the drug set with the largest evaluation disparity for comparison (Fig. 5c).
@@
To further validate the potential of DeepCRE in drug repurposing for CRC, seven more CRC organoids are included for In-SMARchip drug testing. Among the eight CRC organoids tested, five exhibits a significant efficacy increase of set A over set C (Fig. 1d,e, Fig. 6a and Supplementary Fig. 14a). Moreover, drug candidates in set A are also more efficient than two approved drugs in set C in the other three CRC organoids (Supplementary Fig. 14b). Generally, the efficacy patterns of drug candidates were consistent across different patients, except for Sorafenib (D3), one of the receptor tyrosine kinase (RTK) inhibitors (Fig. 6b). However, there were variations in the exact correlation coefficients: inhibitors of PI3K/mTOR signaling and Chromatin histone acetylation show coefficients above 0.96, while those targeting DNA replication display coefficients ranging from 0.33 to 0.93, reflecting the diversity and complexity of chemotherapy [33, 34] (Fig. 3c).
@@
Some intriguing insights emerge from the drug testing results, shedding light on the deeper mechanism exploration of drug candidates with similar MoAs. For instance, the DRC of GSK1059615 exhibits a significant increase at the highest concentration compared to other PI3K/mTOR inhibitors (Fig. 6d), except for C100 (Supplementary Fig. 14c). Perturbation data from CMAP L1000 [35] database provides some explanations (Fig. 6e-f). Cells treated with GSK1059615 show up-regulated cellular survival and stress response related pathways (isomerase activity, intramolecular transferase activity, and phosphotransferases) and genes (IGFBP3, MYCBP, GTF2A2), DNA repair, and cytoskeleton regulation related genes (PARP1, RUVBL1, STAMBP), and down-regulated cell cycle checkpoints related pathways (protein tyrosine/MAP kinase/MAP kinase tyrosine phosphatase activity) and cell cycle related genes (STX1A, PARP6, KDM3A). The up-regulation and down-regulation of these genes and pathways may reflect the defensive mechanisms adopted by cells to resist the toxicity of high concentrations of drugs, potentially leading to the pausing of the cell cycle or the inhibition of cell death pathways [36, 37]. The other example is that the DRC of the drug Sorafenib, rather than another RTK inhibitor Dasatinib, declines at higher concentrations in three CRC organoids (Fig. 6g). It has been reported that some transport proteins may play a role in the accumulation and resistance of drugs, with Dasatinib being primarily restricted by ABCG1, while Sorafenib being mainly limited by ABCG2 [38]. Further investigation suggests that ABCG2 expression, significantly lower in CRC tumor tissues compared to normal tissues [39] (Fig. 6h), may contribute to the reduced resistance of Sorafenib in CRC. Molecular docking analysis corroborates these findings, demonstrating a favorable binding affinity of Sorafenib to the target ABCG2 (Fig. 6i and Supplementary Fig. 14d). Similar conclusions of binding poses and interactions have been drawn by a published paper, which presents single-particle cryo-EM studies of ABCG2 binding with another RTK inhibitor imatinib [40]. Furthermore, survival analysis reveals a negative correlation between ABCG2 expression and disease-free survival (DFS), implicating ABCG2-mediated multi-drug resistance (MDR) in CRC (Fig. 6j).
@@
a, DRC across four CRC organoids, colored by the drug set, indicating that drug set A possesses significantly greater effectiveness than drug set C. b, Heatmap of correlation coefficients of drugs, demonstrating that the drug efficacy is related to the drug MoA and the drug set. c, Boxplot of correlation coefficients of drugs, indicating differences in the distribution of efficient populations among drugs targeting DNA replication. d, DRC of PI3K/mTOR inhibitor drugs across four CRC organoids, showing the dramatic increase in DRC of the drug GSK1059615 at the highest concentration. e, Umap plot of gene expression of cells after different perturbations, with representative differential genes shown in the dashed frame. f, Bar plot of enriched pathways, with up-regulated pathways colored in red and down-regulated pathways in blue. g, DRC of RTK inhibitor drugs across three CRC organoids, showing the decline in DRC of the drug Dasatinib, rather than Sorafenib, at higher concentrations. h, Boxplot of genes ABCG1 and ABCG2 in the TCGA-CRC cohort, indicating significantly lower expression of ABCG2 in tumor samples. i, ABCG2-Sorafenib docking (top) from the view of extracellular space of ABCG2 TM helices in the ATP bound state and (down) in 2D diagram format . j, Survival curve of ABCG2 in the disease-free survival (DFS) of the TCGA-CRC cohort.